OpenStudy (anonymous):

solve the system using substation. Enter answer in ordered pair form (a,b) a+b=4 2a+3b=11

jimthompson5910 (jim_thompson5910):

First isolate b in a+b=4 to get... a+b=4 a+b-a=4-a b = 4-a

jimthompson5910 (jim_thompson5910):

Then plug this into 2a+3b=11 to get 2a+3b=11 2a+3(4-a)=11 ... replace b with 4-a ------------------------------------------------------- solve for 'a' in the equation 2a+3(4-a)=11 gives you a = ???

jimthompson5910 (jim_thompson5910):

what do you get when you distribute 3(4-a)

OpenStudy (anonymous):

12-3

jimthompson5910 (jim_thompson5910):

12-3a actually

jimthompson5910 (jim_thompson5910):

2a+3(4-a)=11 2a+12-3a=11 -a+12 = 11 a = ???

OpenStudy (anonymous):

id subtract12 from 11 correct?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

minus 12 on both sides so you get an equation that looks like -a = -1

jimthompson5910 (jim_thompson5910):

-a = 11 - 12 -a = -1 a = 1

jimthompson5910 (jim_thompson5910):

Once you know a = 1, use it to find that b = 4-a b = 4-1 b = 3

OpenStudy (anonymous):

oh okay o get it.